# Search for Axion dark matter with the QUAX–LNF tunable haloscope

### v1

@@ Abstract
[…]
@@ I Introduction
In recent years, we witnessed an increasing growth in the research of light Dark Matter (DM) candidates, addressing in particular axions and axion-like particles (ALPs). If axions are found to exist, they would untie the long-standing DM problem [1, 2], after being originally postulated as a solution to the strong CP problem [3, 4]. The nature of a pseudoscalar, electrically neutral and feebly interacting particle make the axion a strong DM candidate [5], and its cosmological evolution and astrophysical constraints indicate a favorable mass range between $1~{}\mu\text{eV}<m_{a}<10~{}\text{meV}$ [6].
The research efforts are now spread over many different detection approaches, but the paradigm has become the haloscope design proposed by Sikivie [7, 8], which probes axions from the DM halo of the Galaxy. Currently operating haloscopes are ADMX [9, 10, 11, 12], HAYSTAC [13, 14], ORGAN [15], CAPP-8T [16, 17], CAPP-9T [18], CAPP-PACE [19], CAPP-18T [20], CAST-CAPP [21], GrAHal [22], RADES [23, 24, 25], TASEH [26] and QUAX [27, 28, 29, 30, 31, 32], whereas among the proposed experiments are FLASH [33], ABRACADABRA [34], DM-Radio [35, 36], CADEx [37], MADMAX [38] and ALPHA [39].
@@
$\begin{split}P_{a\gamma}=\left(\frac{g_{a\gamma\gamma}^{2}}{m_{a}^{2}}\,\hbar^{3}c^{3}\rho_{a}\right)\left(\frac{\beta}{1+\beta}\omega_{c}\frac{1}{\mu_{0}}B_{0}^{2}VC_{010}Q_{L}\right)&\\ \times\left(\frac{1}{1+\left(2Q_{L}\,\Delta_{\omega}/\omega_{c}\right)^{2}}\right),\end{split}$ (1)
@@
We assume a local DM density $\rho_{a}=0.45~{}\text{GeV}/\text{cm}^{3}$ [42], $m_{a}$ is the axion mass and $g_{a\gamma\gamma}$ is its coupling constant to photons. $B_{0}$ is the applied magnetic field; $\omega_{c}=2\pi\nu_{c}$, $V$, $Q_{L}$, $\beta$ are respectively the resonance angular frequency of the cavity, the volume, the loaded quality factor and the antenna coupling to the cavity. The relation $Q_{L}=Q_{0}/\left(1+\beta\right)$ holds, with $Q_{0}$ the intrinsic quality factor. $C_{010}$ is a mode dependent geometrical factor, about 0.6 for the TM010 resonant mode of a cylindrical cavity, and $\Delta\omega_{c}=\omega_{c}-\omega_{a}$ is the detuning between the cavity and the axion angular frequency defined as $\omega_{a}=m_{a}c^{2}/\hbar$. The quality factor $Q_{a}\approx 10^{6}$ [43] is related to the energy spread in the cold dark matter halo. In Eq. 2, the signal-to-noise ratio, SNR, is defined as the ratio between the signal power and the uncertainty of the noise power, $\text{SNR}\equiv P_{a\gamma}/\delta P_{\text{noise}}$, while $T_{\text{sys}}$ is the system noise temperature, whose reduction is of fundamental importance in speeding up the scan rate of the experiment.
In this paper, we report on the first operation of the new QUAX haloscope located at the National Laboratories of Frascati (LNF), which doubles the search potential of the QUAX experiment along with the LNL haloscope [32]. The experiment is conducted using a resonant cavity equipped with a tuning rod mechanism allowing to exclude the existence of dark matter axions with coupling $g_{a\gamma\gamma}$ down to $0.861\times 10^{-13}~{}\text{GeV}^{-1}$ in the mass window $(36.5241-36.5510)~{}\mu\text{eV}$.
@@ II.0.1 Haloscope description
The haloscope, schematized in Fig. 1, consists of a cylindrical OFHC copper resonant cavity, with inner radius $r=13.51$ mm and height $h=246$ mm, for a total volume $V=0.141$ l. The body is divided into two semi-cylinders, including the endcaps, sealed together with screws. Two coaxial cables are coupled to the cavity via dipole antennas. One antenna is fixed and weakly coupled, with coupling estimated from simulations $1.4\times 10^{-3}$, such that Eqs. (1) and (2) are still valid. The other is obtained by stripping the end of a coax cable, leaving only the central conductor for a 2 mm length (Bottom-Left panel of Fig. 2). The coax cable is connected to a linear stepper motor whose movement allows the tunability of the coupling $\beta$.
@@
The setup is hosted in a Leiden Cryogenics CF-CS110-1000 dilution refrigerator equipped with two Sumitomo pulse tubes of 1.5 W at 4 K each. The refrigerator is segmented in different temperature stages. The magnet and the HEMT are thermalized at the 4 K stage, while the cavity is connected to the last temperature stage which attained 20 mK at equilibrium. The cavity temperature, monitored during the data taking, reached about 30 mK.
@@ II.0.2 Tuning mechanism
Although, without the tuner, the intrinsic quality factor of the cavity is measured to be $Q_{0}\simeq 10^{5}$ at cryogenic temperatures, the PEEK supporting the rod and non-ideality of the rod itself caused field losses, reducing the quality factor to about 50000 (Sec. II.0.3). The coupling, supposed to be equal to the optimal value for scan $\beta=2$, resulted to be $\beta=0.5$ after a more accurate analysis of calibration data. The values of $Q_{L}$ and $\beta$ measured before the data taking at different frequencies are listed in Tab. 1, while the transmission spectra measured with the VNA are shown in Fig. 3.
@@ II.0.3 Data taking
The data acquisition consists in recording the power coming from the output line L5. During data acquisition, the switch CR7 is closed on line L2 to avoid noise leakage through the under coupled port, the switch RT1 is closed on the line L1 and the switch RT2 on a 50 ohm termination (not shown in Fig. 1). Finally, the switch S1 is closed toward the down conversion and acquisition electronics. Here, with an I–Q mixer we convert the frequency to the baseband, and the I and Q quadratures are amplified by low-noise voltage amplifiers (×$\times 10^{3}$ factor, 10 MHz bandwidth) before being digitized by a 16-bit ADC board, which has a 2 MHz bandwidth and sampling of 2 MS/s. In each sub-run, the I and Q signals are acquired for 4 seconds and saved to file. The total amount of files in each sub-run is 941, resulting in an integration time of $\Delta t=3764$ s.
The calibration procedure is done by a simultaneous fit of the $S_{51}$, $S_{53}$ and $S_{13}$ spectra with their analytical expressions, allowing the extraction of the cavity parameters $\nu_{c}$, $Q_{0}$, $\beta$ and of the attenuation and gain of the input and output lines (see Fig. 4). In particular, we are interested in the gain of Line 5, which is calculated solving a system of three coupled equations:
$S_{51}=L1+s_{21}^{\scriptscriptstyle\text{cav}}(\nu_{c},Q_{0},\beta)+L5,$
$S_{53}=L3+s_{22}^{\scriptscriptstyle\text{cav}}(\nu_{c},Q_{0},\beta)+L5,$ (3)
$S_{13}=L1+s_{12}^{\scriptscriptstyle\text{cav}}(\nu_{c},Q_{0},\beta)+L3,$
where $s_{ij}^{\scriptscriptstyle\text{cav}}$ are the intrinsic transmission and reflection coefficients of the resonant cavity, having Lorentzian shape. The measured gain is $G_{5}\approx 70.0$ dB through all the sub-runs, with a systematic uncertainty of 0.4 dB. This gain is referred at the input of the splitter, having subtracted the cable contributions from the splitter to the VNA and the $-3$ dB of the splitter itself. Separately, the gain and the spectrum shape from the splitter output to the ADC are measured, taking into account the effect of the ADC internal filters.
@@ III Data Analysis and Results
For each frequency step, we estimate $T_{\rm sys}$ from the output power, properly converted into units of Kelvin, at a reference frequency $\nu_{\rm LO}+100~{}\text{kHz}$, where 1/f noise is negligible and still far enough from the cavity resonance, obtaining an average value of 4.7 K. Across the dataset we observe that the power spectra line shapes, show features typical of Fano interference. We interpret this as due to a not perfect isolation and thermalization of the circulator and attenuators, leading to an excess noise temperature of about 600 mK. Similar interference phenomena were recently reported in microwaves measurements [45]. This doesn’t affect the extraction of the power generated by axion conversion since it adds incoherently.
In presence of axion conversions a power surplus is expected in the cavity power spectrum as given by Eq. (1). We calculate the power spectrum residuals by subtracting it to a polynomial obtained with a Savitzky-Golay (SG) filter [46] from the spectrum itself. We adopt a SG filter of the fourth order with a dynamic interval of 250.5 KHz (501 bins). We apply the SG filter in a window [$\nu_{c}-\Gamma,\nu_{c}+\Gamma$] of about 530 kHz, for each dataset. $\Gamma$ is the cavity line-width and is calculated as $\nu_{c}/Q_{L}$. An example of polynomial obtained from a SG filter from one of the datasets is shown in Fig. 5. The slope in the power spectrum is due to aliasing of the thermal noise at the edge of the Nyquist window, while the bump structure is due to the non perfect thermalization and isolation of the rf components connected to the resonant cavity.
@@
where the $\alpha$ index runs over the number of datasets ($N_{\text{scan}}$), the index $i$ runs over the frequency bins of each power spectrum. $S_{\alpha,i}$ and $R_{\alpha,i}$ are the expected power signal and the residuals for frequency bin $i$ and the dataset $\alpha$. We calculate $S_{\alpha,i}$ as the integral in the frequency domain of equation (1) multiplied by the spectrum of the full standard halo model distribution [43].
@@
Minimizing 5 we find: ($\sum\sum\equiv\sum_{\alpha=1}^{N_{\text{scan}}}\sum_{i=1}^{N_{\text{bin}}}$)
@@
To correctly estimate these limits, we estimate the efficiency of the filtering procedure used for the extraction of the axion signal. We follow the procedure reported in [32]. We run a Monte Carlo simulation numerically injecting a fake axion signal, with a known $g_{\text{injected}}^{2}$, into simulated power-spectra with different $\nu_{c}$. We use Eq. (7) to estimate the $g_{\text{injected}}^{2}$, for each injected signal (i.e. obtaining $g_{\text{calculated}}^{2}$). Hence, we determine the efficiency from the relation between $g_{\text{calculated}}^{2}$ and $g_{\text{injected}}^{2}$. To simulate the cavity power-spectra we add random Gaussian noise (with mean=0, sigma=$\sigma_{\scriptscriptstyle\text{Dicke}}$, according to a Gaussian PDF) to the SG filters. Applying this procedure we estimate a detection efficiency of 0.845 on $g_{\text{calculated}}^{2}$ for the SG filter. After correcting for the estimated efficiency, we calculate the limit on the axion-photon coupling with a 90% confidence level as in [31]. We use a power constrained procedure for the $\overline{g^{2}}$ that under fluctuates below $-\sigma$  [48].
Figure 7: The 90% single-sided C.L. upper limit for $g_{a\gamma\gamma}$ as a function of the axion mass. The blue solid curve represents the expected limit in the case of no signal. The yellow region indicates the QCD axion model band. Image realized with [49].
[…]
@@ IV Conclusions
[…]
The apparatus has still much room for improvement. In the upcoming future, the sensitivity can be boosted employing a superconducting resonant cavity, enhancing the intrinsic quality factor. Adding a Josephson Parametric Amplifier as a preamplification stage, which importance have been widely demonstrated in this field, will break the noise temperature down by one order of magnitude. Moreover, the rod design can be optimized to reduce the electromagnetic losses, and substituting the ADC with a 24-bit FPGA would widen the digital dynamics.
To quantify, running such a setup for 1 hour with a quality factor $Q_{0}=3\times 10^{5}$ at a frequency of 9 GHz and coupling $\beta=2$, would reach an average value of $\approx 2\times 10^{-14}~{}\text{GeV}^{-1}$, provided that the noise temperature is 0.5 K and the magnet is brought to 9 T. This value is a factor less than 1.5 from the benchmark KSVZ model, reachable by underfluctuting bins of power.

### v3

@@ Abstract
We report the first experimental results obtained with the new haloscope of the QUAX experiment located at Laboratori Nazionali di Frascati of INFN (LNF). The haloscope is composed of a OFHC Cu resonant cavity cooled down to about 30 mK and immersed in a magnetic field of 8 T. The cavity frequency was varied in a 6 MHz range between 8.831496 and 8.83803 GHz. This corresponds to a previously unprobed mass range between $36.52413$ and $36.5511~{}\mu\text{eV}$. We don’t observe any excess in the power spectrum and set limits on the axion-photon coupling in this mass range down to $g_{a\gamma\gamma}\,\,\textless\,\,0.882\times 10^{-13}$ GeV-1 with the confidence level set to 90%.
@@ I Introduction
In recent years, we witnessed an increasing growth in the research of light Dark Matter (DM) candidates, addressing in particular axions and axion-like particles (ALPs). If axions are found to exist, they would untie the long-standing DM problem [1, 2], after being originally postulated as a solution to the strong CP problem [3, 4]. The nature of a pseudoscalar, electrically neutral and feebly interacting particle makes the axion a strong DM candidate [5, 6, 7], and its cosmological evolution and astrophysical constraints indicate a favorable mass range between $1~{}\mu\text{eV}<m_{a}<10~{}\text{meV}$ [8].
The research efforts are now spread over many different detection approaches, but the paradigm has become the haloscope design proposed by Sikivie [9, 10], which probes axions from the DM halo of the Galaxy. Currently operating haloscopes are ADMX [11, 12, 13, 14], HAYSTAC [15, 16, 17], ORGAN [18, 19], CAPP-8T [20, 21], CAPP-9T [22], CAPP-PACE [23], CAPP-18T [24], CAST-CAPP [25], CAPP-12TB [26], GrAHal [27], RADES [28, 29, 30], TASEH [31] and QUAX [32, 33, 34, 35, 36, 37, 38], whereas among the proposed experiments are FLASH [39], BabyIAXO/RADES [40], ABRACADABRA [41], DM-Radio [42, 43], CADEx [44], MADMAX [45], ALPHA [46], WISPLC [47], DALI [48], BRASS [49], BREAD [50] and SUPAX [51].
@@
$\begin{split}P_{a\gamma}=\left(\frac{g_{a\gamma\gamma}^{2}}{m_{a}^{2}}\,\hbar^{3}c^{3}\rho_{a}\right)\left(\frac{\beta}{1+\beta}\omega_{c}\frac{1}{\mu_{0}}B_{0}^{2}VC_{010}Q_{L}\right)&\\ \times\left(\frac{1}{1+\left(2Q_{L}\,\Delta\omega_{c}/\omega_{c}\right)^{2}}\right),\end{split}$ (1)
@@
We assume a local DM density $\rho_{a}=0.45~{}\text{GeV}/\text{cm}^{3}$ [54], $m_{a}$ is the axion mass and $g_{a\gamma\gamma}$ is its coupling constant to photons. $B_{0}$ is the applied magnetic field; $\omega_{c}=2\pi\nu_{c}$, $V$, $Q_{L}$, $\beta$ are respectively the resonance angular frequency of the cavity, the volume, the loaded quality factor and the antenna coupling to the cavity. The relation $Q_{L}=Q_{0}/\left(1+\beta\right)$ holds, with $Q_{0}$ the intrinsic quality factor. $C_{010}$ is a mode dependent geometrical factor, about 0.69 for the TM010 resonant mode of a cylindrical cavity, and $\Delta\omega_{c}=\omega_{c}-\omega_{a}$ is the detuning between the cavity and the axion angular frequency defined as $\omega_{a}=m_{a}c^{2}/\hbar$. The quality factor $Q_{a}\approx 10^{6}$ [55] is related to the energy spread in the cold dark matter halo. In Eq. (2), the signal-to-noise ratio, SNR, is defined as the ratio between the signal power and the uncertainty of the noise power, $\text{SNR}\equiv P_{a\gamma}/\delta P_{\text{noise}}$, while $T_{\text{sys}}$ is the system noise temperature, whose reduction is of fundamental importance in speeding up the scan rate of the experiment.
In this paper, we report on the first operation of the new QUAX haloscope located at the National Laboratories of Frascati (LNF), which doubles the search potential of the QUAX experiment along with the LNL haloscope [38]. The experiment is conducted using a resonant cavity equipped with a tuning rod mechanism allowing to exclude the existence of dark matter axions with coupling $g_{a\gamma\gamma}$ down to $0.882\times 10^{-13}~{}\text{GeV}^{-1}$ in the mass window $(36.5241-36.5510)~{}\mu\text{eV}$.
@@ II.0.1 Haloscope description
The haloscope, schematized in Fig. 1, consists of a cylindrical OFHC copper resonant cavity, with inner radius $r=13.51$ mm and height $h=246$ mm, for a total volume $V=0.141$ l. The body is divided into two semi-cylinders, including the endcaps, sealed together with screws. Two coaxial cables are coupled to the cavity via dipole antennas. One antenna is fixed and weakly coupled, with coupling estimated from simulations $1.4\times 10^{-3}$ and verified to be less than $7\times 10^{-3}$ from calibration data, such that Eqs. (1) and (2) are still valid. The other is obtained by stripping the end of a coax cable, leaving only the central conductor for a 2 mm length (Bottom-Left panel of Fig. 2). The coax cable is connected to a linear stepper motor whose movement allows the tunability of the coupling $\beta$.
@@
The setup is hosted in a Leiden Cryogenics CF-CS110-1000 dilution refrigerator equipped with two Sumitomo pulse tubes with cooling power of 1.5 W at 4 K each. The refrigerator is segmented into different temperature stages. The magnet and the HEMT are thermalized at the 4 K stage, while the cavity is connected to the last temperature stage which attained 20 mK at equilibrium. The cavity temperature, monitored during the data taking, reached about 30 mK.
@@ II.0.2 Tuning mechanism
Although the intrinsic quality factor of the cavity without the tuner is measured to be $Q_{0}\simeq 10^{5}$ at cryogenic temperatures, the PEEK supporting the rod and non-ideality of the rod itself caused field losses, reducing the quality factor to about 50000 (Sec. II.0.3). The coupling, supposed to be equal to the optimal value for scan $\beta=2$, resulted to be $\beta=0.5$ after a more accurate analysis of calibration data. The values of $Q_{L}$ and $\beta$ measured before the data taking at different frequencies are listed in Tab. 1, while the transmission spectra measured with the VNA are shown in Fig. 3.
@@ II.0.3 Data taking
The data acquisition consists in recording the power coming from the output line L5. During data acquisition, the switch CR7 is closed on line L2 to avoid noise leakage through the under coupled port, the switch RT1 is closed on the line L1 and the switch RT2 is closed on the switch RT3 which, in turn, is closed on a 50 ohm termination (not shown in Fig. 1). Finally, the splitter S1 allows to redirect the signal to the downconversion and acquisition electronics. Here, with an I–Q mixer we convert the frequency to the baseband, and the I and Q quadratures are amplified by low-noise voltage amplifiers (×$\times 10^{3}$ factor, 10 MHz bandwidth) before being digitized by a 16-bit ADC board, which has a 2 MHz bandwidth and sampling of 2 MS/s. In each sub-run, the I and Q signals are acquired for 4 seconds and saved to file. The total amount of files in each sub-run is 941, resulting in an integration time of $\Delta t=3764$ s.
The calibration procedure (refer to Appendix V.2 for more details) is done by a simultaneous fit of the $S_{51}$, $S_{53}$ and $S_{13}$ spectra with their analytical expressions, allowing the extraction of the cavity parameters $\nu_{c}$, $Q_{0}$, $\beta$ and of the attenuation and gain of the input and output lines (see Fig. 4). In particular, we are interested in the gain of Line 5, which is calculated solving a system of three coupled equations (Eq. (10)). The measured gain through all the sub-runs is
$G_{L5}=\left(70.62\pm 0.28_{\scriptscriptstyle\text{syst.}}\pm 0.13_{\scriptscriptstyle\text{stat.}}\right)~{}\text{dB},$ (3)
where the systematic uncertainty is mainly due to the gain flatness of the reconstracted $G_{L5}$, and the statistic uncertainty derives from the data scattering of the $S_{13}$ trace. This gain is referred at the input of the splitter, having subtracted the cable contributions from the splitter to the VNA and the $-3$ dB of the splitter itself. Then, the gain spectrum from the splitter output to the ADC is measured, taking into account the effect of the ADC internal filters.
@@ III Data Analysis and Results
For each frequency step, we estimate $T_{\rm sys}$ from the output power, properly converted into units of Kelvin, at a reference frequency $\nu_{\rm LO}+100~{}\text{kHz}$, where 1/f noise is negligible and still far enough from the cavity resonance, obtaining an average value of 4.7 K. Across the dataset we observe that the power spectra line shapes show features typical of Fano interference, as recently reported in [57] where similar interference phenomena in microwaves measurements were analyzed. We interpret the bump structure as due to a imperfect isolation and thermalization of the circulator, attenuators and the rod inside the cavity, leading to effective temperatures greater than expected from thermalization; in particular we estimate $T_{\textup{circ}}\approx 600$ mK and $T_{\textup{cav}}\approx 400$ mK. This effect does not affect the extraction of the power generated by axion conversion since it adds incoherently.
In presence of axion conversions a power surplus is expected in the cavity power spectrum as given by Eq. (1). We calculate the power spectrum residuals by subtracting it to a polynomial obtained with a Savitzky-Golay (SG) filter [58] from the spectrum itself. We adopt a SG filter of the fourth order with a dynamic interval of 250.5 kHz (501 bins). We apply the SG filter in a window [$\nu_{c}-\Gamma,\nu_{c}+\Gamma$] of about 530 kHz, for each dataset. $\Gamma$ is the cavity line-width and is calculated as $\nu_{c}/Q_{L}$. An example of polynomial obtained from a SG filter from one of the datasets is shown in Fig. 5. The slope in the power spectrum is due to aliasing of the thermal noise at the edge of the Nyquist window, while the bump structure is due to the imperfect thermalization and isolation of the rf components connected to the resonant cavity, as pointed out earlier.
@@
where the $\alpha$ index runs over the number of datasets ($N_{\text{scan}}$), the index $i$ runs over the frequency bins of each power spectrum. $S_{\alpha,i}$ and $R_{\alpha,i}$ are the expected power signal and the residuals for frequency bin $i$ and the dataset $\alpha$. We calculate $S_{\alpha,i}$ as the integral in the frequency domain of Eq. (1) multiplied by the spectrum of the full standard halo model distribution [55].
@@
Minimizing Eq. (5) we find: ($\sum\sum\equiv\sum_{\alpha=1}^{N_{\text{scan}}}\sum_{i=1}^{N_{\text{bin}}}$)
@@
To correctly estimate these limits, we account for the efficiency of the filtering procedure used for the extraction of the axion signal. We follow the procedure reported in [37]. We run a Monte Carlo simulation numerically injecting a fake axion signal, with a known $g_{\text{injected}}^{2}$, into simulated power-spectra with different $\nu_{c}$. We use Eq. (7) to estimate the $g_{\text{injected}}^{2}$, for each injected signal (i.e. obtaining $g_{\text{calculated}}^{2}$). Hence, we determine the efficiency from the relation between $g_{\text{calculated}}^{2}$ and $g_{\text{injected}}^{2}$. To simulate the cavity power-spectra we add random Gaussian noise (with mean=0, sigma=$\sigma_{\scriptscriptstyle\text{Dicke}}$, according to a Gaussian PDF) to the SG filters. Applying this procedure we estimate a detection efficiency of 0.845 on $g_{\text{calculated}}^{2}$ for the SG filter. After correcting for the estimated efficiency, we calculate the single-sided upper limit on the axion-photon coupling with a 90% confidence level as in [36]. We use a power constrained procedure for the $\overline{g^{2}}$ that under fluctuates below $-\sigma$  [60].
Figure 7: Top) The 90% single-sided C.L. upper limit for $g_{a\gamma\gamma}$ as a function of the axion mass. The violet solid curve represents the expected limit in the case of no signal. The yellow region indicates the QCD axion model band. Bottom) Broad view of axion exclusion limits set by haloscopes. The blue vertical line indicates the search presented in this paper. Images realized with [61].
In Fig. 7 we show the upper-limit $g_{a\gamma\gamma}^{\scriptscriptstyle\textup{CL}}$ in an axion mass window of 27.02 neV centered around $36.53764~{}\mu$eV. The maximum sensitivity obtained with a 90% CL is $g_{a\gamma\gamma}^{\scriptscriptstyle\textup{CL}}<0.882\times 10^{-13}$ GeV-1. This value is about 6 times larger with respect to the benchmark QCD axion level of the KSVZ theory [62, 63].
@@ IV Conclusions
In this paper we described the first search for axion dark matter with the tunable QUAX–LNF haloscope. The operation of a microwave cavity resonating at 8.8 GHz and equipped with a movable tuning rod in a magnetic field of 8 T, placed in a stable ultra-cryogenic environment, demonstrated to be successful. Taking data with this setup for a total amount of about 25 hours through 6 MHz allowed us to exclude the existence of axion dark matter in a mass range between $36.52413$ and $36.5511~{}\mu\text{eV}$ with axion-photon coupling $g_{a\gamma\gamma}$ down to $0.882\times 10^{-13}$ GeV-1 with a C.L. of 90%.
The apparatus has still much room for improvement. In the upcoming future, the sensitivity can be boosted employing a superconducting resonant cavity, enhancing the intrinsic quality factor. Adding a Josephson Parametric Amplifier as a preamplification stage, which importance have been widely demonstrated in this field, will break the noise temperature down by one order of magnitude. Moreover, the rod design can be optimized: i. reducing the amount of dielectric exposed to the electromagnetic mode of the cavity, ii. placing the lower end of the bar directly in contact with the endcap of the cavity eliminating PEEK on this side, iii. eventually substituting PEEK with sapphire.
To quantify, running such a setup for 1 hour with a quality factor $Q_{0}=3\times 10^{5}$ at a frequency of 9 GHz and coupling $\beta=2$, would reach an average value of $\approx 2\times 10^{-14}~{}\text{GeV}^{-1}$, provided that the noise temperature is 0.5 K and the magnet is brought to 9 T. This value is a factor less than 1.5 from the benchmark KSVZ model.